What discount rate would make you indifferent between receiving $3,401.00 per year forever and $5,060.00 per year for 29.00 years? Assume the first payment of both cash flow streams occurs in one year.
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=$5060[1-(1+discount rate)^-29]/discount rate

Present value of perpetuity=Annual cash flows/discount rate

=$3401/discount rate

$5060[1-(1+discount rate)^-29]/discount rate =$3401/discount rate

$5060[1-(1+discount rate)^-29] =$3401

[1-(1+discount rate)^-29]=(3401/5060)

1-(3401/5060)=(1+discount rate)^-29

1/(1+discount rate)^29=0.327865612

(1+discount rate)^29=1/0.327865612

(1+discount rate)=(1/0.327865612)^(1/29)

1+Discount rate=1.0392

Discount rate=(1.0392-1)

=3.92%(Approx).